# A Three-Level Whole-Body Disturbance Rejection Control Framework for Dynamic Motions in Legged Robots

### v2

@@ Abstract
This paper presents a control framework designed to enhance the stability and robustness of legged robots in the presence of uncertainties, including model uncertainties, external disturbances, and faults. The framework enables the full-state feedback estimator to estimate and compensate for uncertainties in whole-body dynamics of the legged robots. First, we propose a novel moving horizon extended state observer (MH-ESO) to estimate uncertainties and mitigate noise in legged systems, which can be integrated into the framework for disturbance compensation. Second, we introduce a three-level whole-body disturbance rejection control framework (T-WB-DRC). Unlike the previous two-level approach, this three-level framework considers both the plan based on whole-body dynamics without uncertainties and the plan based on dynamics with uncertainties, significantly improving payload transportation, external disturbance rejection, and fault tolerance. Third, simulations of both humanoid and quadruped robots in the Gazebo simulator demonstrate the effectiveness and versatility of T-WB-DRC. Finally, extensive experimental trials on a quadruped robot validate the robustness and stability of the system when using T-WB-DRC under various disturbance conditions.
@@
This paper presents a practical control framework to significantly improve the robustness of legged robots against real-world uncertainties like unknown payloads, external pushes, and actuator faults. Its core is a novel three-level whole-body controller (T-WB-DRC) that uses a moving horizon estimator (MH-ESO) to accurately identify and compensate for disturbances in real-time. This dual-planning approach, which considers both ideal and disturbance-injected dynamics, outperforms previous methods. The framework’s effectiveness for enhancing stability under disturbance has been successfully validated through extensive simulations and physical experiments on a quadruped robot.
@@ I Introduction
In this paper, we propose a three-level framework that incorporates a full-state feedback estimator and a single processing module at the mid-level to estimate uncertainties in legged systems, addressing model uncertainties, external disturbances, and faults. This framework can include not only the ESO but also other full-state estimators, such as predefined-time observers and model reference adaptive estimators to estimate the uncertainties in legged systems. Furthermore, to mitigate the impact of noise on the legged system, we propose a MH-ESO, which is less sensitive to noise than the ESO, to act as the full-state estimator within the three-level framework. A robust model predictive control, which is less sensitive to noise than QP solving, is adopted to plan the trajectories of dynamics with uncertainties for low-level control. The main contributions and important emphases are the following:
@@ II Preliminaries
To address the uncertainties and noise in legged robots, we propose a three-level whole-body disturbance rejection control architecture, as illustrated in Fig. 2. We employ model predictive control (MPC) in the high level to plan the legged robot’s centroidal momentum, joint positions, joint velocities, and GRFs based on nominal centroidal dynamics. The MH-ESO in middle level estimates disturbances and addresses the noise in the legged system, while the MPC is then used to replan the robot’s joint acceleration and GRFs, based on the reference trajectory from high-level control. Finally, the whole-body control in the low level compensates for system disturbances.
@@ IV Uncertainties Estimation
In this section, we introduce two variants of disturbance estimators, namely ESO and MH-ESO which is designed based on ESO, to estimate uncertainties in the legged system. The disturbance estimators are discussed in Section IV-A. Both disturbance estimators can be adopted in our proposed T-WB-DRC framework. We also present the signal processing module in T-WB-DRC in Section IV-B.
@@ IV-A1 ESO design
where $\hat{u}=[\tau^{\top},F_{\rm{ref}}^{\top}]^{\top}$ with $F_{\rm{ref}}$ being the partial solution of the MPC problem (10) and $\omega_{0}$ is tuning parameter to be determined. We use $F_{\rm{ref}}$ instead of $F$ in (27), because we would like to estimate the discrepancy between $F$ and $F_{\rm{ref}}$ using the ESO.
@@ IV-A2 MH-ESO design
$f_{e}(x_{1},x_{2}):\mathbb{R}^{6+n}\times\mathbb{R}^{6+n}\to\mathbb{R}^{6+n}$ is globally Lipschitz continuous with respect to $x_{1}^{\top},x_{2}^{\top}]^{\top}$. For any two constant vector $a=[a_{1}^{\top},a_{2}^{\top}]^{\top}\in\mathbb{R}^{12+2n}$ and $b=[b_{1}^{\top},b_{2}^{\top}]^{\top}\in\mathbb{R}^{12+2n}$, $f_{e}$ satisfies
@@
$g_{e}(x_{1}):\mathbb{R}^{6+n}\to\mathbb{R}^{6+n}$ is globally Lipschitz continuous with respect to $x_{1}$. For any two constant vector $a_{3},b_{3}\in\mathbb{R}^{6+n}$, $g_{e}$ satisfies
@@
Then, the estimated uncertainties $\hat{f}$ using MH-ESO are given by
@@
In practical applications, to reduce the computation time for solving Problem 1, we do not need to obtain $d_{\rm{opt}}$ using the conclusion in Theorem IV.2 by inverting matrices, but instead can use the sparse LU decomposition method [27] to solve the linear equations (78).
@@ V Robust Model Predictive Control
The derivatives of centroidal momentum are obtained as [20]
@@
where $D_{11}(q)\in\mathbb{R}^{6\times 6}$. The reference acceleration vector for the actuated joints ${\ddot{q}}^{u,j}_{\rm{ref}}\in\mathbb{R}^{n}$ is obtained by directly differentiating the joint velocity $\dot{q}^{u,j}_{\rm{ref}}$. The base reference acceleration $\ddot{q}^{u,b}_{\rm{ref}}$ for the base can be calculate as follows
@@ VII Simulation Verification
To evaluate the effectiveness and versatility of the T-WB-DRC, we conduct a series of simulation experiments on legged robots. First, we present the advantages of the MH-ESO over the ESO. Second, we compare the performance of the proposed T-WB-DRC with the WB-DRC introduced in [13] and the widely adopted standard whole-body control framework (standard WB-C) from [10], while model uncertainties, external disturbances, and fault tolerance are considered.
@@ VII-B Simulations on Quadruped Robots
The operating frequency for high-level control is set to 20 Hz with the prediction horizon $T=1s$ and the total number of steps $H=50$, while the low-level control operates at 1000 Hz. The robust MPC also operates at 20 Hz with the prediction horizon $T_{0}=1s$ and the total number of steps $M=50$. We select MH-ESO as the disturbance estimator to estimate uncertainties and use the saturator as the signal processing module, with the parameter $\alpha=25$ set in the saturator. The operating frequency for MH-ESO is set to 1000 Hz, and the configuration of MH-ESO is provided in Table III. The desired base link (see Fig. 6) height is set to 0.31 m and the desired roll angles of the base link is set to 0 rad. To simulate realistic running environments, we introduced Gaussian random noise $\mathcal{N}(0,0.001)$ to each element of the real position $q^{\rm{real}}$, as shown in (124).
@@ VII-C Simulations on Humanoid Robots
To demonstrate the fault tolerance capability of the humanoid robot, we intentionally reduce the output torque of the right leg’s knee joint (see Fig. 10) by 30%. We select MH-ESO as the disturbance estimator to estimate uncertainties and use the MAF as the signal processing module, with the moving horizon window as 3. Fig. 11 and 12 show the height of the humanoid robot’s base link when the knee joint motor’s output torque is reduced by 30%, with the ESO bandwidth set to $\omega_{0}=100$ and $\omega_{0}=300$, respectively. Table V shows the MSE of the humanoid robot’s base link height relative to the desired base link height. These results illustrate that the T-WB-DRC has the better fault tolerance capability than WB-DRC when the ESO bandwidth $\omega_{0}=100$. As shown in Fig. 12, the robot falls at the start of controller operation due to the increased noise influence as the ESO bandwidth is increased. This demonstrates that the T-WB-DRC has the potential to achieve better locomotion performance than WB-DRC, particularly in the presence of noise.
TABLE V: MSE of of Actual Height Relative to Desired Height
@@
Figure 11: The height of the humanoid robot’s base link when the knee joint motor output torque is cut by 30% with $\omega_{0}=100$.
Figure 12: The height of the humanoid robot’s base link when the knee joint motor output torque is cut by 30% with $\omega_{0}=300$.
@@ VIII Experimental Verification
In this section, we first conduct two comparison real-robot experiments on a single PC (Intel i7-13700KF, 3.4 GHz) with the quadruped A1 robot on flat ground, demonstrating the advantages of the proposed T-WB-DRC over WB-DRC and standard WB-C. We then perform a real-robot experiment on challenging terrain outdoor using a low-performance hardware platform, a microcomputer with an Intel i7-1165G7, 2.8 GHz. First, we reduce the knee joint output torque by 50% to evaluate the fault tolerance of the robot using three different methods. Next, the robot trots in place with a 5 kg load, as shown in Fig. 14, to assess its robustness to model uncertainties. Additionally, to demonstrate the control framework’s ability to handle model uncertainties and external disturbances on rough terrain, we conduct experiments with the robot trotting on such terrain while carrying the 5 kg load, highlighting the T-WB-DRC’s low dependence on hardware. The operating frequency for high-level control is set to 40 Hz with the prediction horizon $T=1s$ and the total number of steps $H=34$, while the low-level control operates at 1000 Hz with the prediction horizon $T_{0}=1s$ and the total number of steps $M=34$. The robust MPC also operates at 40 Hz with the prediction horizon $T=1s$ and discrete interval $\delta_{t}=0.03s$. We choose MH-ESO as the disturbances estimator and the saturator as the signal processing module. The desired base link height is set to 0.31 m and the desired roll angles of the base link is set to 0 rad.
@@
During the time intervals of 8.5s to 13.5s, 18.5s to 23.5s, 28.5s to 33.5s, and 38.5s to 43.5s, the output torque of one knee joint in each of the four legs is intentionally reduced by 50%, with a distinct knee joint subjected to torque reduction in each interval. Fig. 15 shows the height and roll angle of the robot’s base link when the output torque of the four knee joints corresponding to the four legs is reduced by 50% for three different methods. As illustrated in Fig. 15, the robot employing T-WB-DRC demonstrates superior fault tolerance compared to both the robot using WB-DRC and the one utilizing standard WB-C. Fig. 16 illustrates the height and roll angles of the robot’s base link while the robot trots in place, carrying a 5 kg load. As shown in Fig. 16, the height and roll angles of the robot’s base link using T-WB-DRC closely follow the desired trajectory, while the robot using the standard WB-C experiences a failure at nearly 10 seconds. These results indicate that T-WB-DRC demonstrates superior robustness to model uncertainties compared to both WB-DRC and standard WB-C.

### v3

# Note to Practitioners

@@ Abstract
This paper presents a control framework designed to enhance the stability and robustness of legged robots in the presence of uncertainties, including model uncertainties, external disturbances, and faults. The framework enables the full-state feedback estimator to estimate and compensate for uncertainties in the whole-body dynamics of the legged robots. First, we propose a novel moving horizon extended state observer (MH-ESO) to estimate uncertainties and mitigate noise in legged systems, which can be integrated into the framework for disturbance compensation. Second, we introduce a three-level whole-body disturbance rejection control framework (T-WB-DRC). Unlike the previous two-level approach, this three-level framework considers both the plan based on whole-body dynamics without uncertainties and the plan based on dynamics with uncertainties, significantly improving payload transportation, external disturbance rejection, and fault tolerance. Third, simulations of both humanoid and quadruped robots in the Gazebo simulator demonstrate the effectiveness and versatility of T-WB-DRC. Finally, extensive experimental trials on a quadruped robot validate the robustness and stability of the system when using T-WB-DRC under various disturbance conditions.
@@
This paper presents a practical control framework to significantly improve the robustness of legged robots against real-world uncertainties like unknown payloads, external pushes, and actuator faults. Its core is a novel three-level whole-body controller (T-WB-DRC) that uses a moving horizon estimator (MH-ESO) to accurately identify and compensate for disturbances in real-time. This dual-planning approach, which considers both ideal and disturbance-injected dynamics, outperforms previous methods. The framework’s effectiveness in enhancing stability under disturbances has been successfully validated through extensive simulations and physical experiments on a quadruped robot.
@@ I Introduction
In this paper, we propose a three-level framework that incorporates a full-state feedback estimator and a single processing module at the mid-level to estimate uncertainties in legged systems, addressing model uncertainties, external disturbances, and faults. This framework can include not only the ESO but also other full-state estimators, such as predefined-time observers and model reference adaptive estimators to estimate the uncertainties in legged systems. Furthermore, to mitigate the impact of noise on the legged system, we propose an MH-ESO, which is less sensitive to noise than the ESO, to act as the full-state estimator within the three-level framework. A robust model predictive control, which is less sensitive to noise than QP solving, is adopted to plan the trajectories of dynamics with uncertainties for low-level control. The main contributions and important emphases are the following:
@@ II Preliminaries
To address the uncertainties and noise in legged robots, we propose a three-level whole-body disturbance rejection control architecture, as illustrated in Fig. 2. The details of the three-level control framework are further illustrated in Fig. 3. We employ model predictive control (MPC) in the high level to plan the legged robot’s centroidal momentum, joint positions, joint velocities, and GRFs based on nominal centroidal dynamics. The MH-ESO in middle level estimates disturbances and addresses the noise in the legged system, while the MPC is then used to replan the robot’s joint acceleration and GRFs, based on the reference trajectory from high-level control. Finally, the whole-body control in the low level compensates for system disturbances. The proposed framework uses an estimator to estimate and compensate for model uncertainties, external disturbances, and faults. However, it treats these challenges the same without considering their individual effects. Internal model uncertainty requires the estimator to distinguish between parameter drift and true state changes, while external disturbances are easier to estimate if their dynamics differ from the system. Measurement noise creates a trade-off between the observer’s bandwidth and its ability to suppress noise. Sudden faults put the most pressure on the estimator, requiring fast detection and compensation to keep the system stable.
@@ IV Uncertainties Estimation
In this section, we introduce two variants of disturbance estimators, namely ESO and MH-ESO, which are designed based on ESO, to estimate uncertainties in the legged system. The disturbance estimators are discussed in Section IV-A. Both disturbance estimators can be adopted in our proposed T-WB-DRC framework. We also present the signal processing module in T-WB-DRC in Section IV-B.
@@ IV-A1 ESO design
where $\hat{u}=[\tau^{\top},F_{\rm{ref}}^{\top}]^{\top}$ with $F_{\rm{ref}}$ being the partial solution of the MPC problem (10) and $\omega_{0}$ is tuning parameter to be determined. While a large $\omega_{0}$ is necessary for fast disturbance estimation, it can also amplify noise. We use $F_{\rm{ref}}$ instead of $F$ in (27), because we would like to estimate the discrepancy between $F$ and $F_{\rm{ref}}$ using the ESO.
@@ IV-A2 MH-ESO design
$f_{e}(x_{1},x_{2}):\mathbb{R}^{6+n}\times\mathbb{R}^{6+n}\to\mathbb{R}^{6+n}$ is globally Lipschitz continuous with respect to $x_{1}^{\top},x_{2}^{\top}]^{\top}$. For any two constant vectors $a=[a_{1}^{\top},a_{2}^{\top}]^{\top}\in\mathbb{R}^{12+2n}$ and $b=[b_{1}^{\top},b_{2}^{\top}]^{\top}\in\mathbb{R}^{12+2n}$, $f_{e}$ satisfies
@@
$g_{e}(x_{1}):\mathbb{R}^{6+n}\to\mathbb{R}^{6+n}$ is globally Lipschitz continuous with respect to $x_{1}$. For any two constant vectors $a_{3},b_{3}\in\mathbb{R}^{6+n}$, $g_{e}$ satisfies
@@
The estimated uncertainties $\hat{f}$ using MH-ESO are given by
@@
Let $\rho_{r}(A)$ denote the spectral radius of a square matrix A. The MH-ESO is input-to-state stable, and the result in [27] can be used to prove this under the condition that $\rho_{r}(A_{0})<1$.
@@
In practical applications, to reduce the computation time for solving Problem 1, we do not need to obtain $d_{\rm{opt}}$ using the conclusion in Theorem IV.2 by inverting matrices. Instead, we can use the sparse LU decomposition method [28] to solve the linear equations (78).
@@
A shorter horizon $N$ improves computational efficiency and increases uncertainty estimation bandwidth, but it is more sensitive to noise. In contrast, a longer $N$ enhances estimation smoothness and performance, though at the expense of higher computational load and reduced uncertainty estimation bandwidth. The parameter $N$ can be chosen to balance the uncertainty estimation bandwidth and the impact of noise.
@@ V Robust Model Predictive Control
The derivatives of centroidal momentum are obtained from [20]
@@
where $f_{c}(x_{i,k},u_{i,k})$ and the constraint function $g_{i}(x_{i,k},u_{i,k})$ is given in (10). The difference between the MPC problem in (10) and the MPC problem in (110) is that $\hat{d}_{c}$ is introduced in (110), but it is not included in the MPC problem in (10).
@@
where $D_{11}(q)\in\mathbb{R}^{6\times 6}$. The reference acceleration vector for the actuated joints ${\ddot{q}}^{u,j}_{\rm{ref}}\in\mathbb{R}^{n}$ is obtained by directly differentiating the joint velocity $\dot{q}^{u,j}_{\rm{ref}}$. The base reference acceleration $\ddot{q}^{u,b}_{\rm{ref}}$ for the base can be calculated as follows
@@ VII Simulation Verification
To evaluate the effectiveness and versatility of the T-WB-DRC, we conduct a series of simulation experiments on legged robots. First, we present the advantages of the MH-ESO over the ESO. Second, we compare the performance of the proposed T-WB-DRC with the WB-DRC introduced in [13] and the widely adopted standard whole-body control framework (standard WB-C) from [10], while model uncertainties, external disturbances, and fault tolerance are considered. Both the standard WB-C and WB-DRC employ a two-level control framework, comprising high-level and low-level controls, with identical high-level control mechanisms. The primary distinction between the WB-DRC and the standard WB-C lies in the integration of an adaptive ESO within the WB-DRC, which estimates uncertainties in the low-level control, while the standard WB-C does not address such uncertainties. In contrast to both the standard WB-C and WB-DRC, the T-WB-DRC adopts a three-level control framework. It retains the same high-level control as the WB-C and WB-DRC, while incorporating the same low-level control as the WB-DRC.
@@
For the T-WB-DRC, both the nominal and robust MPC operate at specific frequencies and run on separate threads from the WBC. When the MPC and WBC operate simultaneously, the robust MPC uses the last reference trajectory calculated from the most recent nominal MPC solution, while the WBC uses the last reference trajectory calculated from the most recent robust MPC solution. Once the MPC optimization problem is solved, the reference trajectory is updated. We employ a read-write lock to facilitate data transfer between the high-level control and mid-level control, as well as between the mid-level control and the WBC in the low-level control.
@@ VII-B Simulations on Quadruped Robots
The operating frequency for high-level control is set to 20 Hz with the prediction horizon $T=1s$ and the total number of steps $H=50$, while the low-level control operates at 1000 Hz. The robust MPC also operates at 20 Hz with the prediction horizon $T_{0}=1s$ and the total number of steps $M=50$. We define the MPC optimization problem solving time as the duration between the start of the MPC calculation and its completion. Note that the MPC optimization problem solving time needs to be less than 50 ms at a frequency of 20 Hz for the simulation comparison. Each step of the MPC optimization must be solved within 50 ms.
We select MH-ESO as the disturbance estimator to estimate uncertainties and use the saturator as the signal processing module, with the parameter $\alpha=25$ set in the saturator. The operating frequency for MH-ESO is set to 1000 Hz, and the configuration of MH-ESO is provided in Table III. The desired base link (see Fig. 6) height is set to 0.31 m and the desired roll angles of the base link is set to 0 rad. To simulate realistic running environments, we introduced Gaussian random noise $\mathcal{N}(0,0.001)$ to each element of the real position $q^{\rm{real}}$, as shown in (124).
@@
Table V presents the MSE of the humanoid robot’s base link height relative to the desired base link height, considering the robot’s response to model uncertainties over the 10 s to 40 s interval, external disturbances from 0 s to 30 s, and fault tolerance during the 0 s to 30 s interval, as observed in the previous three comparison experiments. The results demonstrate that the robot utilizing the T-WB-DRC method outperforms those using the other two methods.
TABLE V: MSE of Actual Height Relative to Desired Height
MSE
Method Model Uncertainties External Disturbances Faults
WB-C 0.0627 0.0017 3.6961e-04
WB-DRC 0.0103 6.1195e-04 2.4302e-04
T-WB-DRC 0.0025 5.4849e-06 2.7885e-05
@@ VII-C Simulations on Humanoid Robots
To demonstrate the fault tolerance capability of the humanoid robot, we intentionally reduce the output torque of the right leg’s knee joint (see Fig. 10) by 30%. We select MH-ESO as the disturbance estimator to estimate uncertainties and use the MAF as the signal processing module, with the moving horizon window as 3. Fig. 11 and 12 show the height of the humanoid robot’s base link when the knee joint motor’s output torque is reduced by 30%, with the ESO bandwidth set to $\omega_{0}=100$ and $\omega_{0}=300$, respectively. Table VI shows the MSE of the humanoid robot’s base link height relative to the desired base link height. These results illustrate that the T-WB-DRC has the better fault tolerance capability than WB-DRC when the ESO bandwidth $\omega_{0}=100$. As shown in Fig. 12, the robot falls at the start of controller operation due to the increased noise influence as the ESO bandwidth is increased. This demonstrates that the T-WB-DRC has the potential to achieve better locomotion performance than WB-DRC, particularly in the presence of noise.
TABLE VI: MSE of Actual Height Relative to Desired Height
@@
Figure 11: The height of the humanoid robot’s base link when the knee joint motor output torque is reduced by 30% with $\omega_{0}=100$.
Figure 12: The height of the humanoid robot’s base link when the knee joint motor output torque is reduced by 30% with $\omega_{0}=300$.
@@ VIII Experimental Verification
In this section, we first present two real-robot comparison experiments conducted on a single PC (Intel i7-13700KF, 3.4 GHz) using the Unitree A1 quadruped robot. The robot is equipped with joint encoders featuring a 15-bit resolution and an IMU operating at a 1 kHz sampling frequency, all performed on flat ground. These experiments highlight the advantages of the proposed T-WB-DRC over both the WB-DRC and the standard WB-C. We then perform a real-robot experiment on challenging terrain outdoors using a low-performance hardware platform, a microcomputer with an Intel i7-1165G7, 2.8 GHz. First, we reduce the knee joint output torque by 50% to evaluate the fault tolerance of the robot using three different methods. Next, the robot trots in place with a 5 kg load, as shown in Fig. 14, to assess its robustness to model uncertainties. Additionally, to demonstrate the control framework’s ability to handle model uncertainties and external disturbances on rough terrain, we conduct experiments with the robot trotting on such terrain while carrying the 5 kg load, highlighting the T-WB-DRC’s low dependence on hardware. The operating frequency for high-level control is set to 40 Hz with the prediction horizon $T=1$ s and the total number of steps $H=34$, while the low-level control operates at 1000 Hz with the prediction horizon $T_{0}=1$ s and the total number of steps $M=34$. The robust MPC also operates at 40 Hz with the prediction horizon $T=1s$ and discrete interval $\delta_{t}=0.03s$. We choose MH-ESO as the disturbances estimator and the saturator as the signal processing module. The desired base link height is set to 0.31 m and the desired roll angle of the base link is set to 0 rad.
@@
During the time intervals of 8.5s to 13.5s, 18.5s to 23.5s, 28.5s to 33.5s, and 38.5s to 43.5s, the output torque of one knee joint in each of the four legs is intentionally reduced by 50%, with a distinct knee joint subjected to torque reduction in each interval. Fig. 15 shows the height and roll angle of the robot’s base link when the output torque of the four knee joints corresponding to the four legs is reduced by 50% for three different methods. As illustrated in Fig. 15, the robot employing T-WB-DRC demonstrates superior fault tolerance compared to both the robot using WB-DRC and the one utilizing standard WB-C. Fig. 16 illustrates the height and roll angles of the robot’s base link while the robot trots in place, carrying a 5 kg load. As shown in Fig. 16, the height and roll angles of the robot’s base link using T-WB-DRC closely follow the desired trajectory, while the robot using the standard WB-C experiences a failure at nearly 10 seconds. These results indicate that T-WB-DRC demonstrates superior robustness to model uncertainties compared to both WB-DRC and standard WB-C. Table VII presents the MSE of the humanoid robot’s base link height relative to the desired height, considering the robot’s response to model uncertainties over the 5 s to 60 s interval and external disturbances from 5 s to 20 s, as observed in the previous two comparison experiments. The results indicate that the robot employing the T-WB-DRC method outperforms those using the other two methods.
@@
TABLE VII: MSE of Actual Height Relative to Desired Height
MSE
Method Model Uncertainties External Disturbances
WB-C 0.0707 0.0017
WB-DRC 3.4190e-04 3.5434e-04
T-WB-DRC 5.3580e-06 6.4525e-06
@@
Remark VIII.1
Due to the introduction of the uncertainty estimator in WB-DRC, this method demonstrates improved performance over the standard WB-C in terms of disturbance rejection and fault tolerance. However, it also exhibits higher noise sensitivity compared to standard WB-C. By incorporating the MH-ESO, which reduces noise sensitivity relative to the standard ESO, the T-WB-DRC achieves lower noise sensitivity than WB-DRC. As a result, a larger bandwidth can be selected for the T-WB-DRC compared to the WB-DRC, leading to better disturbance rejection and fault tolerance in T-WB-DRC. A summary of the advantages and disadvantages of the three methods is provided in Table VIII. The simulation results in Section VII and experimental results in Section VIII verify the explanations above.
TABLE VIII: The Characteristic Comparison for the Three Methods
Method
Characteristic Standard WB-C WB-DRC T-WB-DRC
Disturbance Rejection Basic Superior to WB-C, but inferior to T-WB-DRC Best
Fault Tolerance Basic Superior to WB-C, but inferior to T-WB-DRC Best
Noise Sensitivity Low High Low